# Similarity-Dissimilarity Loss for Multi-label Supervised Contrastive Learning

### v4

@@ Abstract
Supervised contrastive learning has achieved remarkable success by leveraging label information; however, determining positive samples in multi-label scenarios remains a critical challenge. In multi-label supervised contrastive learning (MSCL), multi-label relations are not yet fully defined, leading to ambiguity in identifying positive samples and formulating contrastive loss functions to construct the representation space. To address these challenges, we: (i) first define five distinct multi-label relations in MSCL to systematically identify positive samples, (ii) introduce a novel Similarity-Dissimilarity Loss that dynamically re-weights samples through computing the similarity and dissimilarity factors between positive samples and given anchors based on multi-label relations, and (iii) further provide theoretical grounded proofs for our method through rigorous mathematical analysis that supports the formulation and effectiveness of the proposed loss function. We conduct the experiments across both image and text modalities, and extend the evaluation to medical domain. The results demonstrate that our method consistently outperforms baselines in a comprehensive evaluation, confirming its effectiveness and robustness. Code is available at: https://github.com/guangminghuang/similarity-dissimilarity-loss.
@@ 1 Introduction
Supervised contrastive learning effectively utilizes label information to yield promising results in single-label scenarios [1, 2, 3]. However, identifying positive samples in multi-label supervised contrastive learning (MSCL) remains a challenge [4]. For example, consider a set of images containing cats and puppies, wherein an anchor image depicts a cat; in the single-label paradigm, positive and negative instances can be unambiguously delineated based on their corresponding taxonomic annotations. Conversely, MSCL introduces inherent classification ambiguity when determining whether an image containing both cats and puppies should be designated as a positive or negative sample in relation to the anchor.
A critical question arises: Should a sample be considered positive when its label set partially overlaps with or exactly matches that of the anchor? Currently, three principal strategies exist for identifying positive samples in multi-label scenarios [4]: (i) ALL considers only samples with an exactly matching label set as positive; (ii) ANY identifies samples with any overlapping class with the anchor as positive, and (iii) MulSupCon [4] conceptually aligns with the ANY approach but treats each label independently, thereby generating multiple distinct positive sets for individual anchors.
However, these methods have inherent limitations, since previous research has overlooked the complicated multi-label relations among samples in MSCL. As illustrated in Figure 1, we introduce five distinct set relations among samples to facilitate a more comprehensive identification of positive sets. The ALL strategy exclusively considers relation $R2$ while disregarding the potential contributions of $R3$, $R4$ and $R5$. Furthermore, long-tailed distributions, when tail samples serve as anchors, the ALL strategy’s requirement for exact label matches significantly impedes these tail anchors from identifying adequate positive samples within a limited batch size, potentially degenerating the method to unsupervised contrastive learning in extreme scenarios [5, 6, 7]. Conversely, both ANY and MulSupCon approaches treat relations $R2$, $R3$, $R4$, and $R5$ identically with equivalent weights in contrastive loss functions, which constitutes a suboptimal approach given the inherent differences among these relations. A detailed mathematical analysis of the limitations for these methods is presented in Section 2 and additional discussion of related work in Appendix F.
@@
To address these issues, we define multi-label relations and introduce a novel contrastive loss function. Our contributions are summarized as follows:
  • To the best of our knowledge, we are the first to define multi-label relations in MSCL, which facilitates the identification of complex relations in multi-label scenarios.
  • We introduce similarity and dissimilarity concepts in multi-label scenarios and propose a novel contrastive loss function, termed Similarity-Dissimilarity Loss, which dynamically re-weights based on the computed similarity and dissimilarity factors between positive samples and anchors, guided by multi-label relations.
  • We establish the theoretical foundations of our approach through rigorous mathematical analysis, demonstrating both the formal derivation, and the upper and lower bounds of the weighting factor.
  • We conduct the experiments across both image and text modalities, and extend the evaluation to medical domain. The results demonstrate that our method consistently outperforms baselines in a comprehensive evaluation, confirming its effectiveness and robustness.
@@ 2.1 Preliminaries
Given a batch of $N$ randomly sample/label pairs, $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1,\dots,N}$, where $\boldsymbol{x}_{i}$ denotes the $i$-th sample and $\boldsymbol{y}_{i}$ its corresponding labels. Here, $\boldsymbol{y}_{i}=\{y_{i}^{(l)}\}_{l=1,\dots,L}$ represents the multi-labels of sample $i$, where $y_{i}^{(l)}$ denotes the $l$-th label of sample $i$ and $L$ is the total number of labels for sample $i$. After data augmentation, the training batch consists of $2N$ pairs, $\{\tilde{\boldsymbol{x}}_{j},\tilde{\boldsymbol{y}}_{j}\}_{j=1,\dots,2N}$, where $\tilde{\boldsymbol{x}}_{2i}$ and $\tilde{\boldsymbol{x}}_{2i-1}$ are two random augmentations of $\boldsymbol{x}_{i}$ ($i=1,\dots,N$) and $\tilde{\boldsymbol{y}}_{2i-1}=\tilde{\boldsymbol{y}}_{2i}=\boldsymbol{y}_{i}$. For brevity, we refer to this collection of $2N$ augmented samples as a "batch" [1].
@@ 2.2 Multi-label Supervised Contrastive Loss
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{\boldsymbol{y}}_{p}=\tilde{\boldsymbol{y}}_{i}\}$ (1)
@@
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{\boldsymbol{y}}_{p}\cap\tilde{\boldsymbol{y}}_{i}\neq\varnothing\}$ (2)
@@
$\mathcal{L}_{i}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (3)
Here, $\tau\in\mathbb{R}^{+}$ represents a positive scalar temperature parameter [5], while $\boldsymbol{z}_{k}=Proj(Enc(\tilde{\boldsymbol{x}}_{k}))\in\mathbb{R}^{D_{P}}$ denotes the projected encoded representation [1].
@@
Zhang and Wu [4] propose an approach that considers each label $\tilde{y}_{i}^{(l)}$ independently, forming multiple positive sets for a given anchor sample $i$. For each label $\tilde{y}_{i}^{(l)}\in\tilde{\boldsymbol{y}}_{i}$, the positive set for the MulSupCon is defined as:
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{y}_{p}^{(l)}\in\tilde{\boldsymbol{y}}_{i}\}$ (5)
For each anchor $i$, the multi-label supervised contrastive loss for MulSupCon is represented as follows [4]:
$\mathcal{L}_{i}^{\text{mul}}=\sum_{\tilde{y}_{p}^{(l)}\in\tilde{\boldsymbol{y}}_{i}}\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (6)
@@
$\mathcal{L}^{\text{mul}}=\frac{1}{\sum_{i}|\tilde{\boldsymbol{y}}_{i}|}\sum_{i\in I}\mathcal{L}_{i}^{\text{mul}}$ (7)
@@ 2.3 Multi-label Relations
In the ALL method, the optimization process aims to align with the mean representation of samples sharing identical label sets [4]. As the example that is demonstrated in Figure 1, for a given anchor $i$, the positive set of ALL is:
@@
In the ALL method, the sample $p_{j}$ in $R2$ is designated as positive sample, while those in relations $R3$, $R4$ and $R5$ are excluded from consideration. Specifically, despite their semantic similarity to anchor $i$ that those overlap labels, the feature representations of samples $p_{j}$ where $j\in{3,4,5}$ are forced away from the anchor in the embedding space, as they are treated as negative examples in the contrastive learning paradigm. Consequently, the restricted size of the positive set $|\mathcal{P}(i)|$ results in a mean representation susceptible to statistical variance. Furthermore, the ALL method may inadvertently treat semantically related samples as negative instances in certain scenarios.
@@
(Vector Similarity Under Label Equivalence). Let $i$ be an anchor and $p$ be any sample in the feature space, where $\tilde{\boldsymbol{y}}_{i},\tilde{\boldsymbol{y}}_{p}\in\mathbb{R}^{d}$ denote their respective label vectors. If $\tilde{\boldsymbol{y}}_{p}=\tilde{\boldsymbol{y}}_{i}$, then under the contrastive learning framework [5], their corresponding projected representations $\boldsymbol{z}_{i},\boldsymbol{z}_{p}\in\mathbb{R}^{m}$ satisfy $\boldsymbol{z}_{i}\simeq\boldsymbol{z}_{p}$.
@@
By applying Lemma 1, the corresponding loss terms in Eq. (3) for samples in different relations exhibit approximate equality:
@@
It is evident that $R2$, $R3$, $R4$ and $R5$ represent fundamentally distinct relations, each characterized by different labels and semantic information. However, the ANY method fails to differentiate these subtle label hierarchies, introducing substantial semantic ambiguity. Moreover, in scenarios where samples predominantly share common classes, the averaging mechanism disproportionately emphasizes these shared classes while diminishing the significance of distinctive features [4].
The MulSupCon method employs a positive sample identification mechanism analogous to ANY, samples $p_{j}$, where $j\in{3,4,5}$ are designated as positive instances. However, MulSupCon distinguishes itself by evaluating each label individually and forming multiple positive sets for a single anchor sample. This approach aggregates positive samples based on the number of overlapping labels between the positive samples and the anchor, thereby expanding the space of positive sets:
@@
Subsequently, the loss for $p_{j}$ in Eq. (6) are as follows by Lemma 1:
@@
For this example (see Figure 1), the MulSupCon successfully discriminates $R3$ and $R4$ from $R2$ and $R5$; however, it fails to establish a distinction between $R2$ and $R5$. This limitation arises primarily because MulSupCon exclusively considers the overlapping regions (Similarity  [^3] The definition of Similarity is introduced in Section 2.4) between anchor $i$ and sample $p$ (i.e., The intersection of sets $\mathcal{S}$ and $\mathcal{T}$), while disregarding the complementary non-intersecting domains (Dissimilarity  [^4] The definition of Dissimilarity is introduced in Section 2.4). That is to say, the similarity between positive samples and anchors is considers, but not yet dissimilarity, which is one of critical information for representation learning in MSCL.
@@ 2.4 Similarity-Dissimilarity Loss
$\mathcal{L}_{i}^{\text{our}}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (13)
@@
$\mathcal{K}_{i,p}^{s}=\frac{|\tilde{\boldsymbol{y}}_{p}^{s}|}{|\tilde{\boldsymbol{y}}_{i}|}=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}$ (14)
@@
$\mathcal{K}_{i,p}^{d}=\frac{1}{1+|\tilde{\boldsymbol{y}}_{p}^{d}|}=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}$ (15)
@@
  • $|\tilde{\boldsymbol{y}}_{i}|=|\mathcal{S}|$ denotes the cardinality of the label space $\tilde{\boldsymbol{y}}_{i}$.
  • $|\tilde{\boldsymbol{y}}_{p}^{s}|=|\mathcal{S}\cap\mathcal{T}|$ measures the cardinality of the intersection of sets $\mathcal{S}$ and $\mathcal{T}$.
  • $|\tilde{\boldsymbol{y}}_{p}^{d}|=|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|$ represents the cardinality of the relative complement with respect to sample $p$.
@@
$|\tilde{\boldsymbol{y}}_{p}^{d}|=|\tilde{\boldsymbol{y}}_{p}|-|\tilde{\boldsymbol{y}}_{p}^{s}|\geq 0$ (16)
where $|\tilde{\boldsymbol{y}}_{p}|$ represents the cardinality of the label space associated with sample $p$.
Specifically, the Similarity-Dissimilarity Loss Loss reduces to Eq. (3), when the following conditions are simultaneously satisfied:
[…]
Accordingly, our proposed loss function constitutes a generalized form of the basic supervised contrastive loss (see Eq. (3)). In particular, Eq. (3) represents a particular case of the Similarity-Dissimilarity Loss. Moreover, our contrastive loss unifies both single-label and multi-label supervised contrastive loss functions within a comprehensive form and paradigm.

### v5

@@ Abstract
Supervised contrastive learning has achieved remarkable success by leveraging label information; however, determining positive samples in multi-label scenarios remains a critical challenge. In multi-label supervised contrastive learning (MSCL), multi-label relations are not yet fully defined, leading to ambiguity in identifying positive samples and formulating contrastive loss functions to construct the representation space. To address these challenges, we: (i) systematically formulate multi-label relations in MSCL, (ii) propose a novel Similarity-Dissimilarity Loss, which dynamically re-weights samples based on similarity and dissimilarity factors, (iii) further provide theoretical grounded proofs for our method through rigorous mathematical analysis that supports the formulation and effectiveness, and (iv) offer a unified form and paradigm for both single-label and multi-label supervised contrastive loss. We conduct experiments on both image and text modalities and further extend the evaluation to the medical domain. The results show that our method consistently outperforms baselines in comprehensive evaluations, demonstrating its effectiveness and robustness. Moreover, the proposed approach achieves state-of-the-art performance on MIMIC-III-Full.
@@ 1 Introduction
Supervised contrastive learning effectively utilizes label information to yield promising results in single-label scenarios (Khosla et al., 2020; Zhang et al., 2022; Lin et al., 2023). However, identifying positive samples in multi-label supervised contrastive learning (MSCL) remains a challenge (Zhang & Wu, 2024). For example, consider a set of images containing cats and puppies, wherein an anchor image depicts a cat; in the single-label paradigm, positive and negative instances can be unambiguously delineated based on their corresponding taxonomic annotations. Conversely, MSCL introduces inherent classification ambiguity when determining whether an image containing both cats and puppies should be designated as a positive or negative sample in relation to the anchor.
A critical question arises: Should a sample be considered positive when its label set partially overlaps with or exactly matches that of the anchor? Currently, three principal strategies [^1] We have discussed related methods in Appendix H exist for identifying positive samples in multi-label scenarios (Zhang & Wu, 2024): (i) ALL considers only samples with an exactly matching label set as positive; (ii) ANY identifies samples with any overlapping class with the anchor as positive, and (iii) MulSupCon (Zhang & Wu, 2024) conceptually aligns with the ANY approach but treats each label independently, thereby generating multiple distinct positive sets for individual anchors.
However, these methods have inherent limitations, since previous research has overlooked the complicated multi-label relations among samples in MSCL. As illustrated in Figure 1, we introduce five distinct set relations among samples to facilitate a more comprehensive identification of positive sets. The ALL method exclusively considers relation $R2$ while disregarding the potential contributions of $R3$, $R4$ and $R5$. Furthermore, long-tailed distributions, when tail samples serve as anchors, the requirement of ALL for exact label matches significantly impedes these tail anchors from identifying adequate positive samples within a limited batch size, potentially degenerating the method to unsupervised contrastive learning in extreme scenarios (Chen et al., 2020; He et al., 2020; Zhang et al., 2023). Conversely, both ANY and MulSupCon approaches treat relations $R2$, $R3$, $R4$, and $R5$ identically with equivalent weights in contrastive loss functions, which constitutes a suboptimal approach given the inherent differences among these relations. A detailed mathematical analysis of the limitations for these methods is presented in Section 2 and additional discussion of related work in Appendix H.
@@
To address the above ambiguities of inter-label relations in MSCL, we define multi-label relations and propose a novel contrastive loss function. It is important to emphasize that this study focuses exclusively on inter-label relations in MSCL, rather than on intra-label or hierarchical relations. Accordingly, the term relation in this paper refers specifically to inter-label relations. Our contributions are summarized as follows:
  • We introduce the notion of multi-label relations in MSCL and propose the Similarity-Dissimilarity Loss, which dynamically re-weights positive pairs based on relations.
  • We establish the theoretical foundations of the loss, including formal derivation, bounds of the weighting factor, and its generalization of supervised contrastive loss to both single- and multi-label settings.
  • We validate our approach on image, text, and medical benchmarks, achieving consistent improvements and state-of-the-art performance on the MIMIC datasets.
@@ 2.1 Preliminaries
Given a batch of $N$ randomly sample/label pairs, $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1,\dots,N}$, where $\bm{x}_{i}$ denotes the $i$-th sample and $\bm{y}_{i}$ its corresponding labels. Here, $\bm{y}_{i}=\{y_{i}^{(l)}\}_{l=1,\dots,L}$ represents the multi-labels of sample $i$, where $y_{i}^{(l)}$ denotes the $l$-th label of sample $i$ and $L$ is the total number of labels for sample $i$. After data augmentation, the training batch consists of $2N$ pairs, $\{\tilde{\bm{x}}_{j},\tilde{\bm{y}}_{j}\}_{j=1,\dots,2N}$, where $\tilde{\bm{x}}_{2i}$ and $\tilde{\bm{x}}_{2i-1}$ are two random augmentations of $\bm{x}_{i}$ ($i=1,\dots,N$) and $\tilde{\bm{y}}_{2i-1}=\tilde{\bm{y}}_{2i}=\bm{y}_{i}$. For brevity, we refer to this collection of $2N$ augmented samples as a ”batch” (Khosla et al., 2020).
@@ 2.2 Multi-label Supervised Contrastive Loss
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{\bm{y}}_{p}=\tilde{\bm{y}}_{i}\}$ (1)
@@
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{\bm{y}}_{p}\cap\tilde{\bm{y}}_{i}\neq\varnothing\}$ (2)
@@
$\mathcal{L}_{i}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\bm{z}_{i}\cdot\bm{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\bm{z}_{i}\cdot\bm{z}_{a}/\tau)}$ (3)
Here, $\tau\in\mathbb{R}^{+}$ represents a positive scalar temperature parameter (Chen et al., 2020), while $\bm{z}_{k}=Proj(Enc(\tilde{\bm{x}}_{k}))\in\mathbb{R}^{D_{P}}$ denotes the projected encoded representation (Khosla et al., 2020).
@@
Zhang & Wu (2024) propose an approach that considers each label $\tilde{y}_{i}^{(l)}$ independently, forming multiple positive sets for a given anchor sample $i$. For each label $\tilde{y}_{i}^{(l)}\in\tilde{\bm{y}}_{i}$, the positive set for the MulSupCon is defined as:
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{y}_{p}^{(l)}\in\tilde{\bm{y}}_{i}\}$ (5)
For each anchor $i$, the multi-label supervised contrastive loss for MulSupCon is represented as follows (Zhang & Wu, 2024):
$\mathcal{L}_{i}^{\text{mul}}=\sum_{\tilde{y}_{p}^{(l)}\in\tilde{\bm{y}}_{i}}\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\bm{z}_{i}\cdot\bm{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\bm{z}_{i}\cdot\bm{z}_{a}/\tau)}$ (6)
@@
$\mathcal{L}^{\text{mul}}=\frac{1}{\sum_{i}|\tilde{\bm{y}}_{i}|}\sum_{i\in I}\mathcal{L}_{i}^{\text{mul}}$ (7)
@@ 2.3 Multi-label Relations
In ALL, the optimization process aims to align with the mean representation of samples sharing identical label sets (Zhang & Wu, 2024). As the example that is demonstrated in Figure 1, for a given anchor $i$, the positive set of ALL is:
@@
In ALL, the sample $p_{j}$ in $R2$ is designated as positive sample, while those in relations $R3$, $R4$ and $R5$ are excluded from consideration. Specifically, despite their semantic similarity to anchor $i$ that those overlap labels, the feature representations of samples $p_{j}$ where $j\in{3,4,5}$ are forced away from the anchor in the embedding space, as they are treated as negative examples in the contrastive learning paradigm. Consequently, the restricted size of the positive set $|\mathcal{P}(i)|$ results in a mean representation susceptible to statistical variance. Furthermore, the ALL method may inadvertently treat semantically related samples as negative instances in certain scenarios.
@@
(Vector Similarity Under Label Equivalence). Let $i$ be an anchor and $p$ be any sample in the feature space, where $\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{p}\in\mathbb{R}^{d}$ denote their respective label vectors. If $\tilde{\bm{y}}_{p}=\tilde{\bm{y}}_{i}$, then under the contrastive learning framework (Chen et al., 2020), their corresponding projected representations $\bm{z}_{i},\bm{z}_{p}\in\mathbb{R}^{m}$ satisfy $\bm{z}_{i}\simeq\bm{z}_{p}$.
@@
By applying Lemma 1, the corresponding loss terms in Equation 3 for samples in different relations exhibit approximate equality:
@@
It is evident that $R2$, $R3$, $R4$ and $R5$ represent fundamentally distinct relations, each characterized by different labels and semantic information. However, ANY fails to differentiate these subtle label hierarchies, introducing substantial semantic ambiguity. Moreover, in scenarios where samples predominantly share common classes, the averaging mechanism disproportionately emphasizes these shared classes while diminishing the significance of distinctive features (Zhang & Wu, 2024).
MulSupCon employs a positive sample identification mechanism analogous to ANY, samples $p_{j}$, where $j\in{3,4,5}$ are designated as positive instances. However, MulSupCon distinguishes itself by evaluating each label individually and forming multiple positive sets for a single anchor sample. This approach aggregates positive samples based on the number of overlapping labels between the positive samples and the anchor, thereby expanding the space of positive sets:
@@
Subsequently, the loss for $p_{j}$ in Equqation 6 are as follows by Lemma 1:
@@
For this example (see Figure 1), MulSupCon successfully discriminates $R3$ and $R4$ from $R2$ and $R5$; however, it fails to establish a distinction between $R2$ and $R5$. This limitation arises primarily because MulSupCon exclusively considers the overlapping regions (Similarity  [^4] The definition of Similarity is introduced in Section 2.4) between anchor $i$ and sample $p$ (i.e., The intersection of sets $\mathcal{S}$ and $\mathcal{T}$), while disregarding the complementary non-intersecting domains (Dissimilarity  [^5] The definition of Dissimilarity is introduced in Section 2.4). That is to say, the similarity between positive samples and anchors is considers, but not yet dissimilarity, which is one of critical information for representation learning in MSCL.
@@ 2.4 Similarity-Dissimilarity Loss
$\mathcal{L}_{i}^{\text{our}}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\exp(\bm{z}_{i}\cdot\bm{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\bm{z}_{i}\cdot\bm{z}_{a}/\tau)}$ (13)
@@
$\mathcal{K}_{i,p}^{s}=\frac{|\tilde{\bm{y}}_{p}^{s}|}{|\tilde{\bm{y}}_{i}|}=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}$ (14)
@@
$\mathcal{K}_{i,p}^{d}=\frac{1}{1+|\tilde{\bm{y}}_{p}^{d}|}=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}$ (15)
@@
  • $|\tilde{\bm{y}}_{i}|=|\mathcal{S}|$ denotes the cardinality of the label space $\tilde{\bm{y}}_{i}$.
  • $|\tilde{\bm{y}}_{p}^{s}|=|\mathcal{S}\cap\mathcal{T}|$ measures the cardinality of the intersection of sets $\mathcal{S}$ and $\mathcal{T}$.
  • $|\tilde{\bm{y}}_{p}^{d}|=|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|$ represents the cardinality of the relative complement with respect to sample $p$.
@@
$|\tilde{\bm{y}}_{p}^{d}|=|\tilde{\bm{y}}_{p}|-|\tilde{\bm{y}}_{p}^{s}|\geq 0$ (16)
where $|\tilde{\bm{y}}_{p}|$ represents the cardinality of the label space associated with sample $p$.
Importantly, a detailed and mathematical explanation of re-weighting mechanism for the Similarity-Dissimilarity Loss is provided in Section 2.7. Additionally, Appendix A includes the scaling hyperparameter temperature $\tau$ (Appendix A.1), and the rationale for selecting the dissimilarity penalty function (Appendix A.2).
2.4.1 A Unified Form of Supervised Contrastive Loss
Specifically, the Similarity-Dissimilarity Loss reduces to Equation 3, when the following conditions are simultaneously satisfied:
$\left\{\begin{aligned} |\tilde{\bm{y}}_{i}|&=|\tilde{\bm{y}}_{p}^{s}|\\ |\tilde{\bm{y}}_{p}^{d}|&=0\end{aligned}\right$. (17)
Accordingly, our proposed loss function constitutes a generalized form of the basic supervised contrastive loss (see Equation 3). In particular, Equation 3 represents a particular case of the Similarity-Dissimilarity Loss. Moreover, our contrastive loss unifies both single-label and multi-label supervised contrastive loss functions within a comprehensive form and paradigm.